# The Riemann Hypothesis is false

### v43

@@
Let: $T\in\mathbb{R}_{>0}$ be arbitrarily large and $X\geq 2$. By Lemma 12.2 of [5], we know that there exists some $X_{0}\in[X,X+1]$ such that $\frac{\zeta^{\prime}}{\zeta}(\sigma+iX_{0})\ll\log^{2}X$ uniformly for $-1\leq\sigma\leq 2$. Thus we can choose $T$ and $H\in[3T,3T+1]$ such that
@@
$\sum_{T\leq\gamma_{T}\leq H}n^{\rho_{T}-1}$ $=\frac{1}{2\pi i}\Bigg{(}\int_{1+\frac{1}{\log n}+iT}^{1+\frac{1}{\log n}+iH}+\int_{1+\frac{1}{\log n}+iH}^{-1+iH}+\int_{-1+iH}^{-1+iT}+\int_{-1+iT}^{1+\frac{1}{\log n}+iT}\Bigg{)}\frac{\zeta^{\prime}}{\zeta}(s)n^{s-1}\mathrm{d}s$ (2)
@@
$S(x):=\sum_{n\leq x}n^{it}\ll T^{\frac{1}{2}}\log^{3}T$ (5)
uniformly for $\frac{T}{\log^{2}T}\leq x\leq T$. Let $f(n)=n^{\sigma-1}.$ Then by Abel summation, it follows that
$\sum_{n\leq x}n^{\sigma-1+it}=S(T)f(T)-S(T/\log^{2}T)f(T/\log^{2}T)-\int_{T/\log^{2}T}^{T}S(x)f^{\prime}(x)\mathrm{d}x\ll T^{\sigma-\frac{1}{2}}\log^{7}T$ (6)
for $-1\leq\sigma\leq 2.$ Combining (1) and (6) gives $\sum_{\frac{T}{\log^{2}T}\leq n\leq T}I_{2}+I_{4}\ll T^{1/2}\log^{7}T.$ By the log-derivative of the functional equation for $\zeta$, we have
@@
for $\sigma\geq\sigma_{0}>0$. Combining (6) and (7) yields $\sum_{\frac{T}{\log^{2}T}\leq n\leq T}I_{3}\ll T^{1/2}\log^{7}T.$ Thus we arrive at
$\sum_{\frac{T}{\log^{2}T}\leq n\leq T}I_{2}+I_{3}+I_{4}\ll T^{1/2}\log^{7}T.$ (8)
@@
$I_{1}$ $=\frac{(T-H)\Lambda(n)}{2\pi n}+\frac{e}{2\pi i}\sum_{m=2,m\neq n}^{\infty}\Bigg{(}\frac{\Lambda(m)n^{iT}}{m^{1+\frac{1}{\log n}+iT}\log(n/m)}-\frac{\Lambda(m)n^{iH}}{m^{1+\frac{1}{\log n}+iH}\log(n/m)}\Bigg{)}.$ (9)
From now on, let $\varepsilon=\varepsilon_{T}=T^{-1/4}$ and $\frac{T}{\log^{2}T}\leq n\leq T$ (unless specified otherwise). For $2\leq m\leq\frac{T}{(1+\varepsilon)\log^{2}T}$ and $m\geq T(1+\varepsilon)$, note that $\frac{1}{|\log(n/m)|}$ is uniformly $\ll 1/\varepsilon=T^{1/4}$. For $m\neq n$, let $g(n)=\frac{1}{m^{1+\frac{1}{\log n}}\log(n/m)}$ hence $g^{\prime}(n)=\frac{m^{-1-\frac{1}{\log n}}(\log(n)\log(m)-\log^{2}(n)-\log^{2}(m))}{n\log^{2}(n)\log^{2}(n/m)}.$ Note that $g^{\prime}(n)$ is strictly negative, hence $|g^{\prime}(n)|=-g^{\prime}(n).$ Thus in similar fashion to (6), we obtain from (5) that
$\sum_{\frac{T}{\log^{2}T}\leq n\leq T}\frac{n^{it}}{m^{1+\frac{1}{\log n}}\log(n/m)}$ $=S(T)g(T)-S(T/\log^{2}T)g(T/\log^{2}T)-\int_{T/\log^{2}T}^{T}S(x)g^{\prime}(x)\mathrm{d}x$ (10)
$\ll\frac{T^{3/4}\log^{5}T}{m^{1+\frac{1}{\log T}}}-T^{1/2}\log^{3}T\int_{T/\log^{2}T}^{T}g^{\prime}(x)\mathrm{d}x$ (11)
@@
uniformly for $2\leq m\leq\frac{T}{(1+\varepsilon)\log^{2}T}$ and $m\geq T(1+\varepsilon).$ For $n\leq\frac{m}{1+\varepsilon}$ and $n\geq m(1+\varepsilon)$, $\frac{1}{|\log(n/m)|}$ is also uniformly $\ll 1/\varepsilon=T^{1/4}$ so it similarly follows from (5) by Abel summation that
$\sum_{\frac{T}{\log^{2}T}\leq n\leq\frac{m}{1+\varepsilon}\leq T}\frac{n^{it}}{m^{1+\frac{1}{\log n}}\log(n/m)}\ll\frac{T^{3/4}\log^{5}T}{m^{1+\frac{1}{\log T}}}$ (13)
@@
$\sum_{\frac{T}{\log^{2}T}\leq m(1+\varepsilon)\leq n\leq T}\frac{n^{it}}{m^{1+\frac{1}{\log n}}\log(n/m)}\ll\frac{T^{3/4}\log^{5}T}{m^{1+\frac{1}{\log T}}}.$ (14)
We now treat the ranges $m+1\leq n\leq m(1+\varepsilon)$ and $\frac{m}{1+\varepsilon}\leq n\leq m-1$, where $\frac{T}{(1+\varepsilon)\log^{2}T}\leq m\leq T(1+\varepsilon)$. For $m+1\leq n\leq m(1+\varepsilon)$, put $n=m+r$ where $11\leq r\leq m\varepsilon-$ 1 thus $\log(n/m)=\frac{r}{m}(1+O(r/m))$ hence $\frac{1}{\log(n/m)}=\frac{m}{r}(1+O(r/m))=\frac{m}{r}+O(1)$ hence
$\sum_{m+1\leq n\leq m(1+\varepsilon)}\frac{\Lambda(m)n^{iT}}{m^{1+\frac{1}{\log n}+iT}\log(n/m)}$ $=\frac{\Lambda(m)}{m^{iT}}\sum_{m+1\leq n\leq m(1+\varepsilon)}\frac{n^{iT}}{m^{1+\frac{1}{\log n}}\log(n/m)}$ (15)
@@
$\sum_{m+1\leq n\leq m(1+\varepsilon)}\frac{\Lambda(m)n^{iT}}{m^{1+\frac{1}{\log n}+iT}\log(n/m)}$ $=\frac{\Lambda(m)}{e}\sum_{1\leq r\leq T^{1/4}}\frac{1}{r}e^{iT\log(1+\frac{r}{m})}+O(\varepsilon\log m).$ (22)
@@
$\sum_{m+1\leq n\leq m(1+\varepsilon)}\frac{\Lambda(m)n^{iT}}{m^{1+\frac{1}{\log n}+iT}\log(n/m)}$ $=\frac{\Lambda(m)}{e}\sum_{1\leq r\leq T^{1/4}}\frac{1}{r}e^{\frac{irT}{m}}+O(\varepsilon\log^{6}T).$ (23)
@@
$\sum_{\frac{T}{(1+\varepsilon)\log^{2}T}\leq m\leq T(1+\varepsilon)}\sum_{m+1\leq n\leq m(1+\varepsilon)}\frac{\Lambda(m)n^{iT}}{m^{1+\frac{1}{\log n}+iT}\log(n/m)}$ $=\sum_{\frac{T}{(1+\varepsilon)\log^{2}T}\leq m\leq T(1+\varepsilon)}\frac{\Lambda(m)}{e}\sum_{1\leq r\leq T^{1/4}}\frac{1}{r}e^{\frac{irT}{m}}$ (24)
@@
For $\frac{m}{1+\varepsilon}\leq n\leq m-1,$ put $n=m-r$ where $1\leq r\leq m(\varepsilon+O(\varepsilon^{2}))$ so that $\log(n/m)=-\frac{r}{m}(1+O(r/m))$ hence $\frac{1}{\log(n/m)}=-\frac{m}{r}(1+O(r/m))=-\frac{m}{r}+O(1)$. Thus by arguing as above, one arrives at
$\sum_{\frac{T}{(1+\varepsilon)\log^{2}T}\leq m\leq T(1+\varepsilon)}\sum_{\frac{m}{1+\varepsilon}\leq n\leq m-1}\frac{\Lambda(m)n^{iT}}{m^{1+\frac{1}{\log n}+iT}\log(n/m)}$ $=-\sum_{\frac{T}{(1+\varepsilon)\log^{2}T}\leq m\leq T(1+\varepsilon)}\frac{\Lambda(m)}{e}\sum_{1\leq r\leq T^{1/4}}\frac{1}{r}e^{-\frac{irT}{m}}$ (26)
@@
$\sum_{\frac{T}{(1+\varepsilon)\log^{2}T}\leq m\leq T(1+\varepsilon)}\Big{(}\sum_{\frac{m}{1+\varepsilon}\leq n\leq m-1}+\sum_{m+1\leq n\leq m(1+\varepsilon)}\Big{)}\frac{\Lambda(m)n^{iT}}{m^{1+\frac{1}{\log n}+iT}\log(n/m)}$ (28)
@@
Let $\{{y\}}$ denote the fractional part of $y$. Recall that if $u$ is not an integer, then $\sum_{n\leq X}\frac{\sin(2\pi nu)}{n}=\pi(\frac{1}{2}-\{{u\}}+O(1/X))$ as $X\rightarrow\infty$. Hence
@@
$\sum_{\frac{T}{(1+\varepsilon)\log^{2}T}\leq m\leq T(1+\varepsilon)}\Big{(}\sum_{\frac{m}{1+\varepsilon}\leq n\leq m-1}+\sum_{m+1\leq n\leq m(1+\varepsilon)}\Big{)}\frac{\Lambda(m)n^{iT}}{m^{1+\frac{1}{\log n}+iT}\log(n/m)}$ (31)
@@
$\sum_{\frac{T}{(1+\varepsilon)\log^{2}T}\leq m\leq T(1+\varepsilon)}\Big{(}\sum_{\frac{m}{1+\varepsilon}\leq n\leq m-1}+\sum_{m+1\leq n\leq m(1+\varepsilon)}\Big{)}\frac{\Lambda(m)n^{iT}}{m^{1+\frac{1}{\log n}+iT}\log(n/m)}$ (38)
@@
$\sum_{\frac{T}{(1+\varepsilon)\log^{2}T}\leq m\leq T(1+\varepsilon)}\Big{(}\sum_{\frac{m}{1+\varepsilon}\leq n\leq m-1}+\sum_{m+1\leq n\leq m(1+\varepsilon)}\Big{)}\frac{\Lambda(m)n^{iH}}{m^{1+\frac{1}{\log n}+iH}\log(n/m)}$ (40)
$=\frac{2\pi i}{e}\Big{(}(B-1)\frac{H}{2\pi}-\frac{H}{2\pi}\log(3/2\pi)+\frac{T}{2}+o(T)\Big{)}$ (41)
$=\frac{2\pi i}{e}\Big{(}(B-1)\frac{3T}{2\pi}-\frac{3T}{2\pi}\log(3/2\pi)+\frac{T}{2}+o(T)\Big{)}$ (42)
since $H=3T+O(1)$. Combining (42) and (39) gives
$\sum_{\frac{T}{(1+\varepsilon)\log^{2}T}\leq m\leq T(1+\varepsilon)}\Big{(}\sum_{\frac{m}{1+\varepsilon}\leq n\leq m-1}+\sum_{m+1\leq n\leq m(1+\varepsilon)}\Big{)}\frac{\Lambda(m)n^{iT}}{m^{1+\frac{1}{\log n}+iT}\log(n/m)}$ (43)
$-\sum_{\frac{T}{(1+\varepsilon)\log^{2}T}\leq m\leq T(1+\varepsilon)}\Big{(}\sum_{\frac{m}{1+\varepsilon}\leq n\leq m-1}+\sum_{m+1\leq n\leq m(1+\varepsilon)}\Big{)}\frac{\Lambda(m)n^{iH}}{m^{1+\frac{1}{\log n}+iH}\log(n/m)}$ (44)
$=\frac{2\pi i}{e}\Big{(}(1-B)\frac{T}{\pi}+\frac{3T}{2\pi}\log 3-\frac{2T}{\pi}\log(2\pi)+o(T)\Big{)}.$ (45)
Summing (9) for $\frac{T}{\log^{2}T}<n\leq T$, changing the order of summation, invoking (45), (14), (13) and (12) yields
$\sum_{\frac{T}{\log^{2}T}<n\leq T}I_{1}$ $=\frac{T-H}{2\pi}\sum_{\frac{T}{\log^{2}T}<n<T}\frac{\Lambda(n)}{n}+(1-B)\frac{T}{\pi}+\frac{3T}{2\pi}\log 3-\frac{2T}{\pi}\log(2\pi)+o(T)$ (46)
$=-\frac{2T}{\pi}\log\log T+(1-B)\frac{T}{\pi}+\frac{3T}{2\pi}\log 3-\frac{2T}{\pi}\log(2\pi)+o(T).$ (47)
@@
$\sum_{T\leq\gamma_{T}\leq H}n^{\rho_{T}-1}=-\frac{T\Lambda(n)}{\pi n}+O((\log T)\log\log 3n)$ (48)
uniformly for $2\leq n\ll T$. Define $N(y):=\sum_{0<\gamma\leq y}1$. Note that $N(y)=(y/2\pi)\log(y/2\pi)-y/2\pi+O(\log y)$ [5, Corollary 14.3]. Thus, $N(H)-N(T)=N(3T)-N(T)+O(\log T)=(T/\pi)\log(T/2\pi)+(T/2\pi)(-2+3\log 3)+O(\log T)$ and $(T/\pi)\sum_{2\leq m\leq\frac{T}{\log^{2}T}}\frac{\Lambda(m)}{m}=(T/\pi)\log T-(2T/\pi)\log\log T-(B/\pi)T+o(T).$ Hence, summing (48) for $1\leq n\leq\frac{T}{\log^{2}T}$ gives
$\sum_{1\leq n\leq\frac{T}{\log^{2}T}}\sum_{T\leq\gamma_{T}\leq H}n^{\rho_{T}-1}$ $=N(H)-N(T)-\frac{T}{\pi}\sum_{2\leq n\leq\frac{T}{\log^{2}T}}\frac{\Lambda(n)}{n}+o(T)$ (49)
$=\frac{2T}{\pi}\log\log T+\frac{T}{2\pi}(3\log 3-2\log(2\pi)+2B-2)+o(T).$ (50)
@@
$\sum_{1\leq n\leq T}\sum_{T\leq\gamma_{T}\leq H}n^{\rho_{T}-1}$ $=\Big{(}\sum_{1\leq n\leq\frac{T}{\log^{2}T}}+\sum_{\frac{T}{\log^{2}T}<n\leq T}\Big{)}\sum_{T\leq\gamma_{T}\leq H}n^{\rho_{T}-1}$ (51)
$\gg T.$ (52)
@@
$\sum_{1\leq n\leq T}\sum_{T\leq\gamma_{T}\leq H}n^{\rho_{T}-1}=\sum_{T\leq\gamma_{T}\leq H}\sum_{1\leq n\leq T}n^{\rho_{T}-1}\ll T^{\Theta}\log T.$ (54)
@@
From the literature on arXiv and elsewhere, it doesn’t appear like this approach on RH has been tried before. However, the idea of applying Gonek’s formula (48) in estimating sums over zeros of zeta in general, is not new. As one can confirm from the literature, it has proven to be a useful tool for such tasks. It is therefore instructive to sum (9) carefully, e.g. by not rushing to naively bound $n^{it}$ by 1 for $n$ not too close to $m$. Doing so will definitely make one lose important information needed for a sufficiently sharp bound for the sums over $n\leq\frac{m}{1+\varepsilon}$ and $n\geq m(1+\varepsilon)$. Note that the analogue of (48) from the Riemann explicit formula is $\sum_{T\leq\gamma_{T}\leq H}\frac{n^{\rho_{T}-1}}{\rho_{T}}\ll\frac{(\log T)(\log\log T)}{T}.$ This has no main term, so it’s not quite useful for our purposes.

### v44

@@
Let: $T\in\mathbb{R}_{>0}$ be arbitrarily large and $X\geq 2$. By Lemma 12.2 of [5], we know that there exists some $X_{0}\in[X,X+1]$ such that $\frac{\zeta^{\prime}}{\zeta}(\sigma+iX_{0})\ll\log^{2}X$ uniformly for $-1\leq\sigma\leq 2$. Thus we can choose $T$ and $H\in[2T,2T+1]$ such that
@@
$\sum_{T\leq\gamma_{T}\leq H}(2n)^{\rho_{T}-1}$ $=\frac{1}{2\pi i}\Bigg{(}\int_{1+\frac{1}{\log 2n}+iT}^{1+\frac{1}{\log 2n}+iH}+\int_{1+\frac{1}{\log 2n}+iH}^{-1+iH}+\int_{-1+iH}^{-1+iT}+\int_{-1+iT}^{1+\frac{1}{\log 2n}+iT}\Bigg{)}\frac{\zeta^{\prime}}{\zeta}(s)(2n)^{s-1}\mathrm{d}s$ (2)
@@
$S(x):=\sum_{n\leq x}(2n)^{it}\ll T^{\frac{1}{2}}\log^{3}T$ (5)
uniformly for $\frac{T}{2\log^{2}T}\leq x\leq T$. Let $f(n)=(2n)^{\sigma-1}.$ Then by Abel summation, it follows that
$\sum_{\frac{T}{\log^{2}T}\leq 2n\leq T}(2n)^{\sigma-1+it}=S(T/2)f(T/2)-S(T/2\log^{2}T)f(T/2\log^{2}T)-\int_{T/2\log^{2}T}^{T/2}S(x)f^{\prime}(x)\mathrm{d}x\ll\frac{\log^{7}T}{T^{\frac{1}{2}-\sigma}}$ (6)
for $-1\leq\sigma\leq 2.$ Combining (1) and (6) gives $\sum_{\frac{T}{\log^{2}T}\leq 2n\leq T}I_{2}+I_{4}\ll T^{1/2}\log^{8}T.$ By the log-derivative of the functional equation for $\zeta$, we have
@@
for $\sigma\geq\sigma_{0}>0$. Combining (6) and (7) yields $\sum_{\frac{T}{\log^{2}T}\leq 2n\leq T}I_{3}\ll T^{1/2}\log^{8}T.$ Thus we arrive at
$\sum_{\frac{T}{\log^{2}T}\leq 2n\leq T}I_{2}+I_{3}+I_{4}\ll T^{1/2}\log^{7}T.$ (8)
@@
$I_{1}$ $=\frac{(T-H)\Lambda(2n)}{4\pi n}+\frac{e}{2\pi i}\sum_{m=2,m\neq 2n}^{\infty}\Bigg{(}\frac{\Lambda(m)(2n)^{iT}}{m^{1+\frac{1}{\log 2n}+iT}\log(2n/m)}-\frac{\Lambda(m)(2n)^{iH}}{m^{1+\frac{1}{\log 2n}+iH}\log(2n/m)}\Bigg{)}.$ (9)
From now on, let $\varepsilon=\varepsilon_{T}=T^{-1/4}$ and $\frac{T}{\log^{2}T}\leq 2n\leq T$ (unless specified otherwise). For $2\leq m\leq\frac{T}{(1+\varepsilon)\log^{2}T}$ and $m\geq T(1+\varepsilon)$, note that $\frac{1}{|\log(2n/m)|}$ is uniformly $\ll 1/\varepsilon=T^{1/4}$. For $m\neq 2n$, let $g(n)=\frac{1}{m^{1+\frac{1}{\log 2n}}\log(2n/m)}$ hence $g^{\prime}(n)=\frac{m^{-1-\frac{1}{\log 2n}}(\log(2n)\log(m)-\log^{2}(2n)-\log^{2}(m))}{n\log^{2}(2n)\log^{2}(2n/m)}.$ Note that $g^{\prime}(n)$ is strictly negative, hence $|g^{\prime}(n)|=-g^{\prime}(n).$ Thus in similar fashion to (6), we obtain from (5) that
$\sum_{\frac{T}{\log^{2}T}\leq 2n\leq T}\frac{(2n)^{it}}{m^{1+\frac{1}{\log 2n}}\log(2n/m)}$ $=S(T/2)g(T/2)-S(T/2\log^{2}T)g(T/2\log^{2}T)-\int_{T/2\log^{2}T}^{T/2}S(x)g^{\prime}(x)\mathrm{d}x$ (10)
$\ll\frac{T^{3/4}\log^{5}T}{m^{1+\frac{1}{\log T}}}-T^{1/2}\log^{3}T\int_{T/2\log^{2}T}^{T/2}g^{\prime}(x)\mathrm{d}x$ (11)
@@
uniformly for $2\leq m\leq\frac{T}{(1+\varepsilon)\log^{2}T}$ and $m\geq T(1+\varepsilon).$ For $2n\leq\frac{m}{1+\varepsilon}$ and $2n\geq m(1+\varepsilon)$, $\frac{1}{|\log(2n/m)|}$ is also uniformly $\ll 1/\varepsilon=T^{1/4}$ so it similarly follows from (5) by Abel summation that
$\sum_{\frac{T}{\log^{2}T}\leq 2n\leq\frac{m}{1+\varepsilon}\leq T}\frac{(2n)^{it}}{m^{1+\frac{1}{\log 2n}}\log(2n/m)}\ll\frac{T^{3/4}\log^{5}T}{m^{1+\frac{1}{\log T}}}$ (13)
@@
$\sum_{\frac{T}{\log^{2}T}\leq m(1+\varepsilon)\leq 2n\leq T}\frac{(2n)^{it}}{m^{1+\frac{1}{\log 2n}}\log(2n/m)}\ll\frac{T^{3/4}\log^{5}T}{m^{1+\frac{1}{\log T}}}.$ (14)
We now treat the ranges $m+1\leq 2n\leq m(1+\varepsilon)$ and $\frac{m}{1+\varepsilon}\leq 2n\leq m-1$, where $\frac{T}{(1+\varepsilon)\log^{2}T}\leq m\leq T(1+\varepsilon)$. For $m+1\leq 2n\leq m(1+\varepsilon)$, put $2n=m+r$ where $11\leq r\leq m\varepsilon-$ 1 thus $\log(2n/m)=\frac{r}{m}(1+O(r/m))$ hence $\frac{1}{\log(2n/m)}=\frac{m}{r}(1+O(r/m))=\frac{m}{r}+O(1)$ hence
$\sum_{m+1\leq 2n\leq m(1+\varepsilon)}\frac{\Lambda(m)(2n)^{iT}}{m^{1+\frac{1}{\log 2n}+iT}\log(2n/m)}$ $=\frac{\Lambda(m)}{m^{iT}}\sum_{m+1\leq 2n\leq m(1+\varepsilon)}\frac{(2n)^{iT}}{m^{1+\frac{1}{\log 2n}}\log(2n/m)}$ (15)
@@
$\sum_{m+1\leq 2n\leq m(1+\varepsilon)}\frac{\Lambda(m)(2n)^{iT}}{m^{1+\frac{1}{\log 2n}+iT}\log(2n/m)}$ $=\frac{\Lambda(m)}{e}\sum_{1\leq r\leq T^{1/4}}\frac{1}{r}e^{iT\log(1+\frac{r}{m})}+O(\varepsilon\log m).$ (22)
@@
$\sum_{m+1\leq 2n\leq m(1+\varepsilon)}\frac{\Lambda(m)(2n)^{iT}}{m^{1+\frac{1}{\log 2n}+iT}\log(2n/m)}$ $=\frac{\Lambda(m)}{e}\sum_{1\leq r\leq T^{1/4}}\frac{1}{r}e^{\frac{irT}{m}}+O(\varepsilon\log^{6}T).$ (23)
@@
$\sum_{\frac{T}{(1+\varepsilon)\log^{2}T}\leq m\leq T(1+\varepsilon)}\sum_{m+1\leq 2n\leq m(1+\varepsilon)}\frac{\Lambda(m)(2n)^{iT}}{m^{1+\frac{1}{\log 2n}+iT}\log(2n/m)}$ $=\sum_{\frac{T}{(1+\varepsilon)\log^{2}T}\leq m\leq T(1+\varepsilon)}\frac{\Lambda(m)}{e}\sum_{1\leq r\leq T^{1/4}}\frac{1}{r}e^{\frac{irT}{m}}$ (24)
@@
For $\frac{m}{1+\varepsilon}\leq 2n\leq m-1,$ put $2n=m-r$ where $1\leq r\leq m(\varepsilon+O(\varepsilon^{2}))$ so that $\log(2n/m)=-\frac{r}{m}(1+O(r/m))$ hence $\frac{1}{\log(2n/m)}=-\frac{m}{r}(1+O(r/m))=-\frac{m}{r}+O(1)$. Thus by arguing as above, one arrives at
$\sum_{\frac{T}{(1+\varepsilon)\log^{2}T}\leq m\leq T(1+\varepsilon)}\sum_{\frac{m}{1+\varepsilon}\leq 2n\leq m-1}\frac{\Lambda(m)(2n)^{iT}}{m^{1+\frac{1}{\log 2n}+iT}\log(2n/m)}$ $=-\sum_{\frac{T}{(1+\varepsilon)\log^{2}T}\leq m\leq T(1+\varepsilon)}\frac{\Lambda(m)}{e}\sum_{1\leq r\leq T^{1/4}}\frac{1}{r}e^{-\frac{irT}{m}}$ (26)
@@
$\sum_{\frac{T}{(1+\varepsilon)\log^{2}T}\leq m\leq T(1+\varepsilon)}\Big{(}\sum_{\frac{m}{1+\varepsilon}\leq 2n\leq m-1}+\sum_{m+1\leq 2n\leq m(1+\varepsilon)}\Big{)}\frac{\Lambda(m)(2n)^{iT}}{m^{1+\frac{1}{\log 2n}+iT}\log(2n/m)}$ (28)
@@
Let $\{{y\}}$ denote the fractional part of $y$. Recall that if $u$ is not an integer, then $\sum_{n\leq X}\frac{\sin(2\pi nu)}{n}=\pi(\frac{1}{2}-\{{u\}}+O(1/X))$ as $X\rightarrow\infty$. Hence if $T\in\mathbb{Q}$ so that $T/2\pi m$ is irrational for $m\in\mathbb{N}$, we have
@@
$\sum_{\frac{T}{(1+\varepsilon)\log^{2}T}\leq m\leq T(1+\varepsilon)}\Big{(}\sum_{\frac{m}{1+\varepsilon}\leq 2n\leq m-1}+\sum_{m+1\leq 2n\leq m(1+\varepsilon)}\Big{)}\frac{\Lambda(m)(2n)^{iT}}{m^{1+\frac{1}{\log 2n}+iT}\log(2n/m)}$ (31)
@@
$\sum_{\frac{T}{(1+\varepsilon)\log^{2}T}\leq m\leq T(1+\varepsilon)}\Big{(}\sum_{\frac{m}{1+\varepsilon}\leq 2n\leq m-1}+\sum_{m+1\leq 2n\leq m(1+\varepsilon)}\Big{)}\frac{\Lambda(m)n^{iT}}{m^{1+\frac{1}{\log n}+iT}\log(2n/m)}$ (38)
@@
$\sum_{\frac{T}{(1+\varepsilon)\log^{2}T}\leq m\leq T(1+\varepsilon)}\Big{(}\sum_{\frac{m}{1+\varepsilon}\leq 2n\leq m-1}+\sum_{m+1\leq 2n\leq m(1+\varepsilon)}\Big{)}\frac{\Lambda(m)(2n)^{iH}}{m^{1+\frac{1}{\log 2n}+iH}\log(2n/m)}$ (40)
$=\frac{2\pi i}{e}\Big{(}(B-1)\frac{H}{2\pi}+\frac{H}{2\pi}\log\pi+\frac{T}{2}+o(T)\Big{)}$ (41)
$=\frac{2\pi i}{e}\Big{(}(B-1)\frac{T}{\pi}+\frac{T}{\pi}\log\pi+\frac{T}{2}+o(T)\Big{)}$ (42)
since $H=2T+O(1)$. Combining (42) and (39) gives
$\sum_{\frac{T}{(1+\varepsilon)\log^{2}T}\leq m\leq T(1+\varepsilon)}\Big{(}\sum_{\frac{m}{1+\varepsilon}\leq 2n\leq m-1}+\sum_{m+1\leq 2n\leq m(1+\varepsilon)}\Big{)}\frac{\Lambda(m)(2n)^{iT}}{m^{1+\frac{1}{\log 2n}+iT}\log(2n/m)}$ (43)
$-\sum_{\frac{T}{(1+\varepsilon)\log^{2}T}\leq m\leq T(1+\varepsilon)}\Big{(}\sum_{\frac{m}{1+\varepsilon}\leq 2n\leq m-1}+\sum_{m+1\leq 2n\leq m(1+\varepsilon)}\Big{)}\frac{\Lambda(m)(2n)^{iH}}{m^{1+\frac{1}{\log 2n}+iH}\log(2n/m)}$ (44)
$=\frac{2\pi i}{e}\Big{(}(1-B)\frac{T}{\pi}+\frac{T}{\pi}\log(\pi/2)+o(T)\Big{)}.$ (45)
Note that $\sum_{m\leq x}\frac{\Lambda(2m)}{2m}=\Big{(}\sum_{m=1}^{\infty}-\sum_{m\geq x+1}\Big{)}\frac{\Lambda(2m)}{2m}=\log 2+O(1/x)$. Therefore, summing (9) for $\frac{T}{\log^{2}T}<2n\leq T$, changing the order of summation, invoking (45), (14), (13) and (12) yields
$\sum_{\frac{T}{\log^{2}T}<2n\leq T}I_{1}$ $=\frac{T-H}{2\pi}\sum_{\frac{T}{\log^{2}T}<2n\leq T}\frac{\Lambda(2n)}{2n}+(1-B)\frac{T}{\pi}+\frac{T}{\pi}\log(\pi/2)+o(T)$ (46)
$=(1-B)\frac{T}{\pi}+\frac{T}{\pi}\log(\pi/2)+o(T)$ (47)
@@
$\sum_{T\leq\gamma_{T}\leq H}(2n)^{\rho_{T}-1}=-\frac{T\Lambda(2n)}{4\pi n}+O((\log T)\log\log 6n)$ (48)
uniformly for $1\leq n\ll T$. Hence, summing (48) for $1\leq n\leq\frac{T}{2\log^{2}T}$ gives
$\sum_{1\leq n\leq\frac{T}{2\log^{2}T}}\sum_{T\leq\gamma_{T}\leq H}(2n)^{\rho_{T}-1}$ $=-\frac{T}{2\pi}\sum_{1\leq n\leq\frac{T}{2\log^{2}T}}\frac{\Lambda(2n)}{2n}+o(T)$ (49)
$=-\frac{T}{2\pi}\log 2+o(T).$ (50)
@@
$\sum_{1\leq n\leq\frac{T}{2}}\sum_{T\leq\gamma_{T}\leq H}(2n)^{\rho_{T}-1}$ $=\Big{(}\sum_{1\leq n\leq\frac{T}{2\log^{2}T}}+\sum_{\frac{T}{2\log^{2}T}<n\leq\frac{T}{2}}\Big{)}\sum_{T\leq\gamma_{T}\leq H}(2n)^{\rho_{T}-1}$ (51)
$=(1-B+\log\pi)\frac{T}{\pi}+\frac{T}{2\pi}\log 2+o(T)$ (52)
@@
$\sum_{1\leq n\leq\frac{T}{2}}\sum_{T\leq\gamma_{T}\leq H}(2n)^{\rho_{T}-1}=\sum_{T\leq\gamma_{T}\leq H}\sum_{1\leq n\leq\frac{T}{2}}(2n)^{\rho_{T}-1}\ll T^{\Theta}\log T.$ (54)
@@
From the literature on arXiv and elsewhere, it doesn’t appear like this approach on RH has been tried before. However, the idea of applying Gonek’s formula (48) in estimating sums over zeros of zeta in general, is not new. As one can confirm from the literature, it has proven to be a useful tool for such tasks. It is therefore instructive to sum (9) carefully, e.g. by not rushing to naively bound $(2n)^{it}$ by 1 for $2n$ not too close to $m$. Doing so will definitely make one lose important information needed for a sufficiently sharp bound for the sums over $2n\leq\frac{m}{1+\varepsilon}$ and $2n\geq m(1+\varepsilon)$. Note that the analogue of (48) from the Riemann explicit formula is $\sum_{T\leq\gamma_{T}\leq H}\frac{(2n)n^{\rho_{T}-1}}{\rho_{T}}\ll\frac{(\log T)(\log\log T)}{T}.$ This has no main term, so it’s not quite useful for our purposes.